x over 2 = 12 over 8 what is x
Answer by stanbon(75887)

x over 2 = 12 over 8
-------------------
x/2 = 12/8
---
x = 2(12/8)
---
x = 24/8
----
x = 3
============
Cheers,
Stan H.
====================